### v1

# EvoXplain: When Machine Learning Models Agree on Predictions but Disagree on Why Measuring Mechanistic Multiplicity Across Training Runs

Abstract
Machine learning models are primarily judged by their predictive performance, especially in applied settings. Once a model reaches a high accuracy, its explanation is often assumed to be both correct and trustworthy. However, this assumption raises an often overlooked question: when two models achieve high accuracy, do they depend on the same internal logic, or do they reach the same outcome via different—and potentially competing—mechanisms?
We introduce EvoXplain, a simple diagnostic framework that measures the stability of model explanations across repeated training. Rather than analysing the explanation of a single trained model, EvoXplain treats explanations as samples drawn from the stochastic optimisation process itself—without aggregating predictions or constructing ensembles—and examines whether these samples form a single coherent explanation or instead separate into multiple, distinct explanatory modes.
We evaluate EvoXplain on the Breast Cancer and COMPAS datasets across two widely used model classes: Logistic Regression and tree-based models (Random Forests). Although all models achieve high predictive accuracy, their explanations frequently exhibit clear multimodality. Even models commonly assumed to be stable—such as Logistic Regression—can produce multiple, well-separated explanatory basins under repeated training on the same data split. Importantly, these differences are not explained by hyperparameter variation or simple performance trade-offs.
EvoXplain does not attempt to select a “correct” explanation. Instead, it makes explanatory instability visible and quantifiable, revealing when averaged or single-instance explanations obscure the existence of multiple underlying mechanisms. This provides a practical check on interpretability claims and highlights a failure mode of common explanation-averaging practices used in ensemble and AutoML pipelines. More broadly, EvoXplain reframes interpretability as a property of a model class under repeated instantiation, rather than of any single trained model.
1 Introduction
Machine learning models are increasingly deployed in domains where explanations matter, including healthcare, public policy, and scientific discovery. Despite the fact that practitioners routinely retrain models during development and performance tuning, interpretability is typically assessed using a single trained instance—or by averaging explanations across runs—rather than by explicitly examining whether explanations themselves remain consistent under retraining. This practice implicitly assumes that high predictive accuracy corresponds to a single, stable way of reasoning (lipton2018myth; rudin2019stop).
There have been sustained challenges to this assumption in the literature. Perhaps the most well-known is the Rashomon effect, which observes that distinct models may achieve similar predictive performance while relying on different internal logics (breiman2001statistical; marx2020predictive). Related work on underspecification highlights that modern learning pipelines often admit many solutions that are indistinguishable in terms of accuracy, leaving optimisation procedures to select arbitrarily among them (d2022underspecification). Other studies have shown that explanations can vary substantially when models are retrained, or when features are correlated or perturbed, including documented failures of attribution methods under retraining, feature overlap, or input transformations (kindermans2017reliability; nogueira2018stability; kumar2020problems; frye2020missingness). Collectively, these results suggest that explanation stability cannot be assumed a priori. However, existing work does not provide a practical way to quantify how much explanatory variability exists in a given setting, nor whether such variability is structured or merely diffuse.
To address this gap, we propose EvoXplain, a simple and easy-to-apply framework for assessing the stability of model explanations across repeated training. Rather than evaluating a single attribution vector, EvoXplain treats explanations as outcomes of stochastic optimisation and analyses them as a distribution induced by the training process itself. By clustering explanation samples obtained from repeated instantiations of the same model class—without aggregating predictions or constructing ensembles—EvoXplain determines whether training consistently yields a single explanatory mode or instead produces multiple, distinct explanatory regimes.
EvoXplain does not aim to decide which explanation is correct. Instead, it provides evidence about whether a model’s explanations are reliable or whether they depend sensitively on the particular optimisation trajectory taken during training. This perspective aligns with a growing view that interpretability should be evaluated empirically rather than assumed, a position increasingly reflected in discussions of responsible AI, model auditing, and governance (arrieta2020explainable; selbst2018intuitive; ecegai2023aiact). By making explanation stability measurable, EvoXplain enables practical questions such as: does this model tell a consistent story, or does its reasoning change across retraining?
In this paper, we introduce the EvoXplain framework, demonstrate its application on high-stakes datasets, and discuss why explanation stability matters for scientific inference, auditing, and model governance. Our goal is to support interpretability practice by providing researchers and practitioners with a diagnostic tool for detecting when explanations are stable—and when multiple explanatory stories coexist despite similar predictive performance.
2 Related Work
2.1 Model Performance and Interpretability
A large proportion of studies in machine learning assess models primarily through the lens of accuracy. In the real world, interpretability is often categorized based upon the model being used. Linear models and decision trees, for instance, are often tagged as “interpretable” because practitioners are able to read and comprehend the parameters or structures involved. However, some recent contributions to the field argue that readability cannot be equated with the models achieving high predictive accuracy (lipton2018myth). EvoXplain follows the same direction by considering interpretability as something that can be measured and verified rather than assumed.
2.2 Rashomon Effects and Predictive Multiplicity
Our work relates to the Rashomon effect in that different models can arrive at the same answer and yet rely on totally different paths or reasoning (breiman2001statistical; semengue2019rashomon). Some recent research considers predictive multiplicity under settings of fairness and of unconstrained risk (umbach2023predictive; marx2020predictive; fisher2019all), showing that the same risk scores can be justified through different models (umbach2023predictive). Unlike these approaches, EvoXplain studies multiplicity within the same model class, trained on the same data, to see whether retraining leads to different explanation patterns.
2.3 Underspecification and Training Variability
Research on underspecification d2022underspecification shows that training pipelines may converge to many possible solutions that perform equally well. EvoXplain complements this work by looking specifically at explanation behaviour. Instead of only asking whether predictions differ, EvoXplain checks whether internal reasoning differs even when predictions stay the same.
2.4 Feature Importance Instability
Previous studies indicate that SHAP, LIME, and gradients as explanation methods can produce instability due to retraining, dataset perturbations, or correlated features (kindermans2017reliability; nogueira2018stability; kumar2020problems; agarwal2021uncertainty). Most of these studies focus on instability in some specific values. EvoXplain extends local noise analysis by detecting whether instability forms structured groups of explanations and how often each group appears.
2.5 Stability-Aware Interpretability
Certain approaches quantify uncertainty in explanations, such as reporting ranges or using measures of variance (molnar2022interpretable; agarwal2021uncertainty; visani2022statistical). These methods help describe the extent of dispersion across runs, but usually do not assess whether the dispersion pertains to different explanatory modes. EvoXplain works on identifying and quantifying such modes.
2.6 Mechanistic Interpretability in Deep Models
The understanding of circuits and the internal functioning of large neural networks remains an ongoing area of research (olah2020circuits). Specific techniques tend to rely on handcrafted analyses or other tools engineered for individual designs. On the other hand, EvoXplain focuses on using generic attribution outputs to determine mechanism-level differences.
2.7 Summary
There is an acknowledgment of the following from the previous literature:
  1. 1.
    Different distinguishable models are able to achieve the same level of performance.
  2. 2.
    Optimization may lead the model to have multiple valid solutions.
  3. 3.
    Explanations may differ during retraining.
What is absent is a simple way to quantify explanation variability, when the latter is present. More specifically, we need to understand explanation variability, its diversity, and its frequency. EvoXplain addresses this limitation by turning explanation stability to a visible property instead of an assumption.
3 Problem Formulation
We use interpretability methods to understand how models make decisions, but most analyses implicitly assume that a model has a single, stable explanation once trained. This assumption is not always justified. Due to randomness in the training process, model inductive biases, and model flexibility, it is possible for multiple different explanations to achieve similar predictive performance. A first step toward trustworthy interpretability is therefore to verify whether explanatory multiplicity exists at all.
Formally, consider a model class $\mathcal{A}$ trained on a fixed dataset $\mathcal{D}$. Standard interpretability workflows examine a single trained instance $\theta$ and its corresponding attribution vector $\Phi(\theta)$. In contrast, this work asks the following question: when $\mathcal{A}$ is trained multiple times with independent random initializations, do the resulting attribution vectors converge to a single explanatory pattern, or do they separate into multiple distinct patterns?
Let ${\theta_{r}}_{r=1}^{R}$ denote $R$ independent training runs on $\mathcal{D}$, and let $\mathbf{e}_{r}=\Phi(\theta_{r})$ denote the corresponding explanation vectors. EvoXplain analyses the empirical distribution of $\mathbf{e}_{r}$} to determine:
  1. 1.
    whether explanations concentrate around a single dominant mode,
  2. 2.
    whether they disperse into multiple stable explanatory modes, and
  3. 3.
    how much support each mode receives.
In practice, we repeat this analysis across multiple train–test splits of $\mathcal{D}$ to assess the robustness of explanatory behaviour under data resampling. However, the core object of interest remains the distribution of explanations induced by repeated training within a fixed data realisation.
EvoXplain offers an empirical method for answering these questions. Rather than assuming explanations are stable, it treats them as random samples generated by the training process and assesses their structure using clustering and entropy. This reframes interpretability as an empirical question: whether a model class tends to produce a single coherent explanation or multiple competing explanatory mechanisms under repeated instantiation.
Contributions
This work makes three main contributions.
First, we consider that explanation stability is something to be quantified. Instead of assuming that a model has a single reasoning way, we think of explanations as samples that can be analysed over several training runs.
Second, we introduce EvoXplain, a simple diagnostic framework that determines if explanations cluster into a single pattern or if they are dispersed across multiple clusters of different patterns. It also measures the occurrence frequency of each distinct explanatory pattern.
Third we apply EvoXplain across well-known and studied datasets using numerous widely used model classes. We observe that internal reasoning is not consistent even at high performance accuracy and this shows the need for explanation stability as a measure of model evaluation and supports interpretability approaches that examine the variability of explanatory stories a model class can produce.
4 Methods
EvoXplain treats explanations as outcomes that may vary across training runs, rather than as fixed properties of a single trained model or an average across multiple trained instances. Instead of assuming that a model class converges to a single mode of reasoning, EvoXplain empirically evaluates whether explanations are stable, whether multiple distinct explanatory patterns emerge, and how frequently each pattern occurs. The method follows a simple workflow: repeatedly train a model under fixed conditions, extract explanations, represent them in a common space, and analyse their stability.
4.1 Repeated Training and Attribution Sampling
We initially conducted the EvoXplain analysis using Gini-based feature importance and observed the same qualitative phenomenon of structured explanatory multiplicity across repeated training runs; we focus on SHAP in this paper because it is the attribution method most commonly used in deployed models and high-stakes decision-making contexts.
Let $\mathcal{A}$ denote a model class (e.g., Logistic Regression or Random Forests). For each run $r=1,\dots,R$, we train $\mathcal{A}$ from an independent random seed while keeping the data split, preprocessing, and hyperparameters fixed. After each training run, we compute feature-level attributions using a chosen interpretability method (here, SHAP), yielding an attribution vector $\mathbf{e}_{r}\in\mathbb{R}^{d}$ over the $d$ input features.
EvoXplain collects these attribution vectors,
$E=\{\mathbf{e}_{r}\}_{r=1}^{R},$
and treats them as samples drawn from the explanation distribution induced by stochastic training. Importantly, EvoXplain does not aggregate predictions or construct ensembles; it analyses explanation variability arising solely from repeated instantiation of the same model class.
4.2 Representing Explanation Space
To enable comparison across runs, attribution vectors are normalised to remove scale effects and focus on relative feature importance. Specifically, we centre each vector using the mean attribution across runs and apply $\ell_{2}$ normalisation:
$\hat{\mathbf{e}}_{r}=\frac{\mathbf{e}_{r}-\mu}{\|\mathbf{e}_{r}-\mu\|_{2}},$
where $\mu$ is the mean attribution across runs. The resulting matrix $\hat{E}\in\mathbb{R}^{R\times d}$ represents the set of explanations as points in a shared geometric space.
4.3 Clustering and Basin Structure
To assess whether explanations converge to a single pattern or separate into multiple patterns, we apply $k$-means clustering (chosen for simplicity and interpretability) for values of $k$ from $1$ to $K_{\max}$. For each $k$, we compute the silhouette score $S(k)$ and select
$k^{\star}=\arg\max_{k}S(k).$
If $k^{\star}=1$, explanations concentrate around a single explanatory mode. If $k^{\star}>1$, EvoXplain identifies multiple explanatory basins corresponding to distinct explanation patterns reached across training runs.
4.4 Mechanistic Entropy and Basin Support
Let $c_{r}\in{1,\dots,k^{\star}}$ denote the cluster assignment of run $r$. The empirical support of basin $j$ is defined as
$p_{j}=\frac{1}{R}\sum_{r=1}^{R}\mathbb{I}[c_{r}=j].$
We define mechanistic entropy as the Shannon entropy of this distribution,
$H=-\sum_{j=1}^{k^{\star}}p_{j}\log(p_{j}),$
normalised by $\log(k^{\star})$ (with $H=0$ if $k^{\star}=1$), so that $H\in[0,1]$. Lower values indicate concentration around a dominant basin, while values near one indicate that training runs are broadly distributed across multiple explanatory basins.
4.5 Centroid Profiles and Mode Interpretation
Each explanatory basin $j$ is associated with a centroid
$\bar{\mathbf{e}}_{j}=\frac{1}{|\{r:c_{r}=j\}|}\sum_{r:c_{r}=j}\hat{\mathbf{e}}_{r},$
which represents the characteristic explanation pattern of that basin. Comparing centroid profiles reveals how and where the model’s reasoning diverges across runs, for example by emphasising different features or feature combinations.
4.6 Mechanistic Diagnosis
EvoXplain provides three core diagnostics:
  1. 1.
    Mode count: whether explanations converge ($k^{\star}=1$) or split into multiple modes.
  2. 2.
    Mechanistic entropy: how training runs are distributed across explanatory modes.
  3. 3.
    Basin profiles: the characteristic feature importance patterns of each mode.
These diagnostics do not identify a “correct” explanation; interpretation is left to domain experts and regulatory context.
4.7 Algorithm Summary
EvoXplain
1:Train $\mathcal{A}$ for $R$ independent seeds; collect attribution vectors $E={\mathbf{e}_{r}}$
2:Normalise: $\hat{\mathbf{e}}_{r}\leftarrow(\mathbf{e}_{r}-\mu)/|\mathbf{e}_{r}-\mu|1$
3:for $k=1$ to $K{\max}$ do
4:  Cluster explanations into $k$ clusters
5:  Compute silhouette score $S(k)$
6:end for
[…]
8:Assign each run to cluster $c_{r}\in{1,\dots,k^{\star}}$
9:Compute basin supports $p_{j}$ and entropy $H$
10:Compute centroid profiles $\bar{\mathbf{e}}_{j}$
11:return $(k^{\star},H,{p_{j}},{\bar{\mathbf{e}}_{j}})$
Clustering in explanation space is used here as a diagnostic tool rather than as a claim about the true number of underlying mechanisms. We adopt k-means with silhouette-based selection for simplicity and transparency, and because the resulting clusters are visually apparent in low-dimensional projections and consistent across independent train–test splits. Importantly, the presence or absence of explanation multiplicity does not hinge on fine-grained cluster boundaries: when training converges to a single explanatory regime (e.g., under fixed regularisation), explanations collapse to a single dense region and silhouette analysis consistently selects k = 1, whereas when multiple explanatory regimes are admissible, well-separated modes emerge with stable support across runs and splits. EvoXplain therefore does not rely on clustering to create structure, but uses clustering to quantify structure that is already present in the explanation distribution.
4.8 Theoretical Interpretation
EvoXplain can be understood through optimisation non-identifiability. Let $\Theta$ denote the parameter space of a model class and $\mathcal{L}(\theta)$ its loss surface. In high-dimensional or weakly constrained settings, training may converge to multiple parameter configurations with near-optimal loss,
$\mathcal{M}=\{\theta\in\Theta:\mathcal{L}(\theta)\approx\mathcal{L}^{\ast}\},$
which we refer to as a mechanism manifold. Different training runs may select different points in $\mathcal{M}$ and yield distinct attribution patterns even when predictive performance is similar.
Let $\Phi:\Theta\rightarrow\mathbb{R}^{d}$ map model parameters to attribution vectors. EvoXplain studies the empirical distribution
$\mathcal{E}=\{\Phi(\theta_{r}):\theta_{r}\sim\text{Train}(\mathcal{A})\}.$
where clustering reveals groups of trained model instances with similar explanatory behaviour. Mechanistic entropy quantifies how training disperses across these groups. When $k^{\star}=1$, training concentrates near a single attractor; when $k^{\star}>1$, training repeatedly reaches different regions of the explanation manifold.
EvoXplain does not claim these basins correspond to true causal mechanisms. Rather, it demonstrates that optimisation and model capacity alone can induce multiple, stable explanatory behaviours.
4.9 Computational Considerations
For $R$ training runs and clustering over $C$ candidate values of $k$, the computational cost scales as $O(RT+RCd)$, where $T$ is the training time for one model and $d$ the number of features. Since training and attribution computations are independent across runs, EvoXplain is naturally parallelisable. In practice, a few hundred runs are often sufficient to reveal stable explanatory structure, while larger values of $R$ improve resolution.
4.10 Attribution Method Considerations
[…]
4.11 Diagnostic Applications
EvoXplain can serve as a screening tool for interpretability audits, fairness assessments, scientific analysis, and model governance. By identifying cases where explanations fail to stabilise, it signals when explanation use or aggregation should be treated with caution. In high-stakes domains such as healthcare, criminal risk assessment, and public policy, EvoXplain can highlight when additional scrutiny or domain-specific review is warranted.
5 Scope and Limitations
EvoXplain does not assert that the clusters it identifies correspond to true causal mechanisms, nor that particular centroids represent any form of ground truth. Its outputs are contingent upon the attribution method used, and the observed structure may also reflect noise or inductive biases inherent to the model or the explanation technique. EvoXplain does not attempt to resolve these questions. Rather, its purpose is to bring such issues to the forefront and to encourage more cautious interpretation of explanation outputs.
EvoXplain has been designed to assess how models behave across repeated training runs. It determines whether a model settles into a single cluster of explanations or whether different explanations emerge across retraining. This perspective comes with several limitations.
First, while this work applies EvoXplain using attribution-based explanations, the framework itself is more general. Attribution vectors represent one way of operationalising interpretability, but other classes of explanations—such as causal abstractions, counterfactuals, symbolic rules, or circuit-level features—may also be suitable, provided they yield representations that can be quantified across multiple runs. As such, our findings instantiate $\Theta$ using a particular family of explanation paradigms; extending EvoXplain to other forms of explanation is a natural direction for future work.
Second, EvoXplain detects the presence of multiple explanatory basins but does not assign any of them the status of a “correct” basin. Mechanism selection, validation, or prioritisation requires domain knowledge, causal evidence, and, in many settings, regulatory judgement. EvoXplain is intended to support such processes, not replace them.
Third, our empirical analysis is limited to tabular datasets and a small number of widely used model classes. While these settings are well studied and practically relevant, EvoXplain has not yet been evaluated across all data modalities, model architectures, or explanation techniques.
Fourth, EvoXplain assumes that repeated training is feasible. For many tabular and classical machine learning models this assumption is easily satisfied. For large-scale models, such as foundation systems, full retraining may be impractical, and alternative strategies—such as checkpoint sampling or multiple fine-tuning runs—may be required.
These limitations do not diminish the framework, but rather clarify the scope of its current applicability. EvoXplain is a diagnostic tool: it does not provide definitive answers, but instead surfaces stability issues that are often overlooked. Future work may extend the framework to additional explanation paradigms, model classes, data types, causal validation strategies, and decision criteria for interpreting and resolving basin structure.
6 Experimental Setup
6.1 Datasets
We evaluate EvoXplain on two tabular binary classification datasets.
  • Breast Cancer Wisconsin (Diagnostic): $n=569$, $d=30$. The task is to classify tumors as malignant or benign. Data are loaded from sklearn.datasets.load_breast_cancer.
  • COMPAS Recidivism: $n\approx 7,200$. The task is to predict two-year recidivism. We use the ProPublica COMPAS dataset with standard filtering (screening window, valid charge degree, non-missing scores). Categorical variables are one-hot encoded.
For each dataset, we perform stratified train–test splits with test fraction $0.3$ and fixed split seeds. All analyses are repeated independently per split.
6.2 Models
We evaluate the following model classes:
  • Logistic Regression (LogReg): $\ell_{2}$-regularised logistic regression with solvers lbfgs (deterministic).
  • Random Forest (RF): Bootstrap aggregation with varying tree hyperparameters.
Linear models are trained on standardised features using StandardScaler. Tree-based models are trained on raw features.
6.3 Repeated Training Protocol
For each of the 10 dataset–model–split combinations (with five splits used for Random Forests), we perform $R=1000$ independent training runs.
Each run differs only in:
  • the random seed used for optimisation and internal sampling,
  • the regularisation or model hyperparameters when explicitly varied.
For Logistic Regression, the regularisation strength $C$ is either:
  • fixed ($C=1.0$), or
  • varied deterministically over a log-spaced grid between $10^{-2}$ and $10^{2}$.
For Random Forests, hyperparameters are sampled from predefined grids over: number of estimators, maximum depth, minimum samples per leaf, and maximum features.
Each training run produces a trained model, test accuracy, and a global explanation vector. Runs are executed in parallel on HPC infrastructure using chunked execution and aggregated post hoc (See the code repository). Due to the computational cost of Random Forest explanations, RF experiments were conducted over five independent train–test splits, which were sufficient to characterise the qualitative geometry of the explanation landscape.
Although the lbfgs solver is deterministic for fixed data and fixed hyperparameters, explanation multiplicity arises because training spans a range of regularisation strengths, allowing optimisation to settle into different equally optimal basins across the hyperparameter-conditioned loss landscape. Thus, the stochasticity measured by EvoXplain reflects basin selection under admissible training configurations, not randomness in optimisation or attribution.
6.4 Explanation Generation
For each trained model, we compute global feature importance using SHAP.
  • Logistic Regression: shap.LinearExplainer
  • Random Forest: shap.TreeExplainer
SHAP values are computed on the held-out test set. For each run $r$, we obtain a single global importance vector
$\mathbf{e}_{r}\in\mathbb{R}^{d}$
by averaging absolute SHAP values across test samples and outputs.
6.5 Aggregation and Clustering
For each split, explanation vectors are aggregated across runs and reordered using run indices to ensure completeness.
Explanation vectors are $\ell_{2}$-normalised prior to clustering. Degenerate cases (nearly identical vectors) are detected using a fixed tolerance.
Clustering is performed using $k$-means with candidate values
$k\in\{2,3,4,5,6,7,8\}.$
The optimal number of clusters $k^{\star}$ is selected by maximising the silhouette score. Cluster centroids, sizes, and mechanistic entropy are computed as described in 4.3.
6.6 Disagreement Inspection
When $k^{\star}>1$, we identify representative runs for each explanatory basin by selecting the run closest to the cluster centroid in cosine distance.
Representative models are retrained and evaluated on the same test set. For each test instance $x$, we compute the disagreement score
$\Delta(x)=\max_{i}P(y{=}1\mid x,M_{i})-\min_{i}P(y{=}1\mid x,M_{i}),$
where $M_{i}$ denotes a representative model from basin $i$.
Instances with the largest $\Delta(x)$ are reported to characterise cases where distinct explanatory modes lead to materially different predictions.
6.7 Visualisation
For each split, explanation vectors are projected into two dimensions using PCA after centring. Per-split manifolds display cluster structure and regularisation gradients.
A universal manifold is constructed by stacking explanation vectors across all splits and applying PCA without per-split centring, enabling comparison of global geometric structure across data realisations.
7 Results
7.1 Logistic Regression: explanation multiplicity under stochastic training
7.1.1 Breast Cancer
We start with Logistic Regression on the Breast Cancer dataset, since this is the simplest model in our study and, in principle, one where we would expect explanations to be stable. When the regularisation strength is allowed to vary, this expectation does not hold.
For a representative split (split seed 104), we analyse 1,000 independent training runs. In explanation space, the SHAP vectors do not form a single cloud. Instead, they organise along a smooth, low-dimensional manifold with clear separation between regions (Fig. 1). Clustering in explanation space yields an optimal solution at k=2, with a silhouette score of 0.608 (compared to 0.566 for k=3 and steadily decreasing thereafter), indicating well-separated explanation modes. The normalized explanation entropy is 0.991 (maximum possible value 1.0), showing that runs are almost evenly distributed across the two modes rather than collapsing into one dominant explanation.
[…]
These differences are reflected in the explanations themselves. In split 104, the dominant features in Cluster 0 are classical size-related measurements such as radius error, worst radius, and worst area, whereas Cluster 1 places substantially more weight on error-based and shape-related features such as fractal dimension error and area error, with absolute SHAP magnitudes several times larger than those seen in Cluster 0. Thus, the model is not merely reweighing the same features slightly differently, but relying on qualitatively different explanatory patterns.
Importantly, the average attribution vector—computed as the arithmetic mean of all SHAP vectors across runs—does not lie on the populated manifold. In the split-level embedding, it falls in a low-density region between clusters (Fig. 1), indicating that this average explanation does not correspond to a typical trained model instance.
When we stack explanation vectors across all splits, the same qualitative picture persists. The universal explanation manifold shows multiple distinct tracks that remain separated across resampled datasets, while predictive accuracy remains high along all tracks (Fig. 2). The global average attribution vector, computed across all runs and all clusters, again lies away from the high-density regions of the manifold, reinforcing that averaging explanations across stochastic runs produces a vector that is not representative of deployed models.
To assess whether these explanation differences can translate into different predictions at the individual level, we compare representative models from the two clusters in split 104. Across held-out test instances, the mean absolute disagreement in predicted malignant probability is 0.049, but the tail of the distribution is substantial, with a maximum disagreement of 0.848. For example, for test instance 30 (ground truth: benign), one representative model assigns a malignant probability of 0.15, while the other assigns 0.999. Similar large disagreements are observed for multiple benign and malignant cases, with disagreement scores exceeding 0.5 for several individuals. These discrepancies arise despite both models achieving high overall accuracy.
Control Experiment: Stability under Fixed Regularization.
As a control, we repeat the same analysis under fixed regularisation strength ($C=1.0$). Across ten independent train-test splits, each with 1000 runs, the explanation vectors consistently collapse to a single cluster ($k=1$), with negligible dispersion in explanation space. This confirms that the observed explanation multiplicity is not an artefact of the SHAP method or dimensionality reduction, but emerges specifically when the training process admits multiple equally accurate models.
Figure 1: Breast Cancer — Logistic Regression. Left: PCA embedding of explanation (attribution) vectors obtained from repeated training runs on the same train–test split. Points are coloured by cluster assignment identified in explanation space. Red stars indicate representative cluster centroids, while the black cross denotes the average attribution vector across all runs. Right: The same embedding with predictive accuracy overlaid. Accuracy remains high across the full extent of the explanation manifold, including across distinct clusters, indicating that multiple explanatory modes coexist at comparable performance.
Figure 2: Breast Cancer - Logistic Regression: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking runs across all train-test splits. Each point corresponds to a trained model instance and is coloured by predictive accuracy. Multiple persistent explanation tracks are visible across splits, indicating that explanation variability is not confined to a single resampling of the data. The black $\times$ marks the global average attribution vector, computed as the arithmetic mean of all SHAP vectors across runs and clusters; this point does not correspond to a typical location on the populated manifold.
Table 1: Split-level summary for Logistic Regression on Breast Cancer under stochastic training. For each train-test split, we report the number of runs, the selected number of explanation clusters $k^{*}$, the silhouette score at $k^{*}$, the normalized cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). This table quantifies split-to-split consistency of the clustering outcome and confirms that multiplicity is not driven by unstable performance.
Split Runs $k^{*}$ Silh. $(k^{*})$ $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 3 0.554 0.972 0.973 0.004
101 1000 2 0.603 1.000 0.960 0.008
102 1000 3 0.540 0.982 0.971 0.010
103 1000 2 0.585 1.000 0.977 0.012
104 1000 2 0.608 0.991 0.967 0.010
105 1000 2 0.604 0.985 0.956 0.005
106 1000 2 0.647 0.997 0.967 0.009
107 1000 3 0.562 0.983 0.958 0.005
108 1000 2 0.616 0.992 0.969 0.009
109 1000 2 0.603 0.984 0.963 0.007
Table 2: Cluster-level summary for Logistic Regression on Breast Cancer under stochastic training. Each row corresponds to one explanation cluster within one train–test split. We report cluster size, regularisation strength statistics (mean $\pm$ std and range of $C$ values across runs assigned to the cluster), and predictive accuracy (mean $\pm$ std). This table characterises how explanatory basins align with distinct regularisation regimes while maintaining similar predictive performance.
Split Cl. Runs $C$ mean $C$ std $C_{\min},C_{\max}]$ Acc. mean Acc. std
100 0 293 3.275 2.418 [0.639, 9.4] 0.974 0.003
100 1 451 0.151 0.161 [0.010, 0.6] 0.974 0.005
100 2 256 38.547 25.163 [9.528, 100.0] 0.971 0.002
101 0 504 0.221 0.259 [0.010, 1.0] 0.964 0.009
101 1 496 21.742 25.113 [1.042, 100.0] 0.957 0.004
102 0 275 36.484 25.431 [7.997, 100.0] 0.963 0.004
102 1 429 0.129 0.132 [0.010, 0.5] 0.972 0.011
102 2 296 2.726 2.031 [0.522, 7.9] 0.977 0.007
103 0 498 0.212 0.245 [0.010, 1.0] 0.976 0.015
103 1 502 21.494 25.064 [0.986, 100.0] 0.979 0.006
104 0 556 0.325 0.410 [0.010, 1.7] 0.971 0.012
104 1 444 24.132 25.496 [1.684, 100.0] 0.963 0.006
105 0 573 0.369 0.477 [0.010, 2.0] 0.956 0.005
105 1 427 25.021 25.599 [1.969, 100.0] 0.955 0.004
106 0 532 0.272 0.332 [0.010, 1.3] 0.971 0.007
106 1 468 22.972 25.330 [1.349, 100.0] 0.962 0.008
107 0 283 35.670 25.519 [7.428, 100.0] 0.953 0.000
107 1 427 0.127 0.130 [0.010, 0.5] 0.961 0.005
107 2 290 2.575 1.883 [0.513, 7.4] 0.957 0.003
108 0 449 23.882 25.463 [1.608, 100.0] 0.968 0.005
108 1 551 0.313 0.392 [0.010, 1.6] 0.969 0.011
109 0 574 0.372 0.481 [0.010, 2.0] 0.965 0.005
109 1 426 25.075 25.604 [1.987, 100.0] 0.959 0.008
[…]
At the same time, predictive performance remains stable. Mean accuracy across splits ranges from 0.956 to 0.977, with standard deviations typically below 0.012. There is no split in which increased explanation multiplicity coincides with degraded or unstable accuracy, consistent with the split-level manifolds where distinct explanation modes occupy similar performance bands.
[…]
Clusters are not small or degenerate. Cluster sizes are substantial and often roughly balanced (for example, split 104 yields clusters of size 556 and 444). Predictive accuracy remains similar across clusters within each split, with differences typically on the order of a few percentage points or less and overlapping accuracy distributions throughout.
Therefore, Tables 2 and 1 show that explanation multiplicity in Logistic Regression on Breast Cancer is consistent across splits, involves large fractions of runs, and aligns with distinct regularisation regimes, while predictive performance remains uniformly high. As a control, when regularisation strength is fixed, explanations collapse to a single cluster (k=1) across all splits and runs (See code repository), confirming that the observed multiplicity arises specifically when the training process admits multiple equally accurate solutions.
7.1.2 COMPAS
Results for Logistic Regression on the COMPAS dataset are summarised in Figures 3 and 4 and in Tables 3 and 4. Across all ten train–test splits, explanation clustering consistently selects two clusters (k=2), with silhouette scores tightly concentrated between 0.82 and 0.83 (Table 3). These values indicate strong and stable separation between explanation modes across resampled datasets. In contrast to the Breast Cancer dataset, the normalized entropy is consistently lower (approximately 0.56-0.58), indicating that explanation modes are unequally populated rather than evenly balanced.
Predictive performance remains stable across splits. Mean accuracy ranges from 0.67 to 0.69, with standard deviations on the order of $10^{-3}$ or smaller (Table 3). There is no evidence that the emergence of multiple explanation modes is associated with unstable or degraded performance. This is consistent with the split-level manifolds (Fig.3), where accuracy varies smoothly along the explanation geometry rather than separating sharply by cluster.
[…]
When explanation vectors are stacked across all splits, the same structure persists (Fig. 4). The universal explanation manifold consists of multiple separated tracks that remain distinct across resampled datasets, while accuracy remains confined to a narrow band. As in the split-level analysis, the global average attribution vector lies away from the populated regions of the manifold, indicating that averaging explanations across stochastic runs does not yield a representative explanation.
As a control, fixing the regularisation strength results in a collapse to a single explanation cluster across all splits (refer to the results folder in the provided code repository). Therefore, the COMPAS results show that explanation multiplicity in Logistic Regression is highly consistent across splits and strongly tied to distinct regularisation regimes, while predictive performance remains effectively unchanged.
Figure 3: COMPAS — Logistic Regression (single split). Left: PCA embedding of explanation (SHAP) vectors obtained from repeated training runs on a fixed train–test split (split 105). Points are coloured by explanation cluster assignment. Red stars indicate representative cluster centroids, and the black $\times$ denotes the average attribution vector across all runs. Right: The same embedding with predictive accuracy overlaid. Accuracy remains within a narrow range across the explanation manifold, indicating that distinct explanatory modes coexist at similar predictive performance.
Figure 4: COMPAS - Logistic Regression: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking model runs across all train-test splits. Each point corresponds to a trained model instance and is coloured by predictive accuracy. Multiple separated explanation tracks persist across splits, indicating that explanation variability is not confined to a single resampling of the data. The black $\times$ marks the global average attribution vector, computed as the arithmetic mean of all SHAP vectors across runs and clusters; this point lies away from the populated tracks of the manifold.
Table 3: Split-level summary for Logistic Regression on COMPAS under stochastic training. For each train-test split, we report the number of runs, the selected number of explanation clusters $k^{*}$, the silhouette score at $k^{*}$, the normalized cluster entropy $H_{\mathrm{norm}}$, and predictive accuracy (mean $\pm$ std across runs). This table quantifies split-to-split consistency of the clustering outcome and confirms that multiplicity is not driven by unstable performance.
Split Runs $k^{*}$ Silh. $(k^{*})$ $H_{\mathrm{norm}}$ Acc. mean Acc. std
100 1000 2 0.829 0.568 0.683 0.001
101 1000 2 0.830 0.568 0.685 0.001
102 1000 2 0.830 0.566 0.694 0.001
103 1000 2 0.822 0.574 0.671 0.001
104 1000 2 0.826 0.566 0.681 0.001
105 1000 2 0.824 0.579 0.679 0.001
106 1000 2 0.830 0.566 0.676 0.000
107 1000 2 0.830 0.568 0.678 0.003
108 1000 2 0.830 0.574 0.674 0.002
109 1000 2 0.829 0.571 0.686 0.001
Table 4: Cluster-level summary for Logistic Regression on COMPAS under stochastic training. Each row corresponds to one explanation cluster within one train-test split. We report cluster size, regularisation strength statistics (mean $\pm$ std and range of $C$ values across runs assigned to the cluster), and predictive accuracy (mean $\pm$ std). This table characterises how explanatory basins align with distinct regularisation regimes while maintaining similar predictive performance.
Split Cl. Runs $C$ mean $C$ std $C_{\min},C_{\max}]$ Acc. mean Acc. std
100 0 866 12.578 21.767 [0.034, 100.0] 0.683 0.000
100 1 134 0.020 0.007 [0.010, 0.0] 0.682 0.001
101 0 866 12.578 21.767 [0.034, 100.0] 0.685 0.000
101 1 134 0.020 0.007 [0.010, 0.0] 0.684 0.002
102 0 867 12.564 21.759 [0.034, 100.0] 0.694 0.000
102 1 133 0.020 0.007 [0.010, 0.0] 0.692 0.001
103 0 864 12.607 21.784 [0.035, 100.0] 0.671 0.001
103 1 136 0.020 0.007 [0.010, 0.0] 0.671 0.001
104 0 867 12.564 21.759 [0.034, 100.0] 0.682 0.000
104 1 133 0.020 0.007 [0.010, 0.0] 0.681 0.002
105 0 862 12.637 21.801 [0.036, 100.0] 0.679 0.001
105 1 138 0.020 0.007 [0.010, 0.0] 0.676 0.001
106 0 867 12.564 21.759 [0.034, 100.0] 0.676 0.000
106 1 133 0.020 0.007 [0.010, 0.0] 0.676 0.001
107 0 866 12.578 21.767 [0.034, 100.0] 0.679 0.001
107 1 134 0.020 0.007 [0.010, 0.0] 0.672 0.003
108 0 864 12.607 21.784 [0.035, 100.0] 0.674 0.001
108 1 136 0.020 0.007 [0.010, 0.0] 0.670 0.002
109 0 865 12.593 21.776 [0.035, 100.0] 0.687 0.001
109 1 135 0.020 0.007 [0.010, 0.0] 0.683 0.001
7.1.3 Disagreement between equally accurate basins
To assess whether explanation multiplicity can translate into materially different predictions at the individual level, we compare representative models drawn from distinct explanation basins within the same train–test split. For each held-out test instance $i$, we compute the absolute difference in predicted class probability,
$\Delta_{i}=\left|P(y=1\mid x_{i},M_{b_{1}})-P(y=1\mid x_{i},M_{b_{2}})\right|,$ (1)
where $M_{b_{1}}$ and $M_{b_{2}}$ denote representative models from two different explanation clusters.
On the Breast Cancer dataset, disagreement magnitudes span a wide range. Across held-out test instances, the mean disagreement is modest, but the upper tail is substantial. In split 104, the maximum disagreement reaches $\Delta=0.85$, with several cases exceeding $\Delta=0.4$. These large discrepancies occur despite both models achieving similarly high overall accuracy. The highest-disagreement cases correspond to specific individuals for whom one model assigns a near-certain malignant (or benign) probability while another assigns a substantially lower risk, indicating that explanation multiplicity can lead to sharply divergent predictions at the individual level.
On the COMPAS dataset, disagreement magnitudes are smaller but non-zero. For split 105, the mean disagreement across test instances is approximately 0.01, with a maximum disagreement of 0.05. Although these values are lower than those observed for Breast Cancer, they still indicate that equally accurate models can assign different recidivism probabilities to the same individual, even when explanation clusters are strongly separated and predictive accuracy is tightly constrained.
These results show that explanation multiplicity is not merely a geometric property of explanation space, but can translate into different model outputs for individual test instances. In settings where multiple explanation basins coexist at similar accuracy, the choice of model instance implicitly determines which individuals receive higher or lower predicted risk, even when aggregate performance metrics are indistinguishable. These clusters therefore represent qualitatively distinct explanatory mechanisms rather than perturbations of a single solution. Despite near-identical test accuracy, the two basins induce materially different prediction behaviour at the individual level, with a maximum disagreement of 0.848. This demonstrates that high predictive performance alone does not imply mechanistic stability or uniqueness.
[…]
To understand the source of these disagreements, we examine the characteristic explanation patterns associated with each basin. Figure 5 visualises the centroid attribution profiles of the two explanatory modes identified on the Breast Cancer dataset, revealing how equally accurate models rely on qualitatively different feature groupings when making predictions.
7.2 Random Forests: explanation multiplicity in ensemble models
Having established that explanation multiplicity arises in Logistic Regression under stochastic training—despite its convex objective and stable predictive performance—we next examine whether the same phenomenon appears in a qualitatively different model class. Random Forests introduce non-convexity, ensemble averaging, and additional sources of stochasticity through bootstrap sampling and feature subsampling. Evaluating explanation stability in this setting allows us to test whether the observed multiplicity is specific to linear models under hyperparameter variation, or whether it persists in more complex and widely deployed learning architectures.
7.2.1 Breast Cancer
Results for Random Forests on the Breast Cancer dataset are shown in Fig. 6 and Table 5; Fig. 7 illustrates that the same multi-modal structure is already present within a single train-test split. Each train-test split aggregates approximately 192,000 Random Forest instances, reflecting stochasticity arising from bootstrap sampling, feature subsampling, and hyperparameter variation. In explanation space, SHAP vectors form multiple dense and well-separated regions rather than the smooth manifolds observed for Logistic Regression. These regions persist across splits, indicating stable explanation multiplicity in the ensemble setting.
Table 5 shows that two explanation clusters are consistently identified in each split, with a stable size ratio of approximately 1:2. Predictive performance remains comparable across clusters: mean accuracy differences are small (typically below 0.01), and standard deviations overlap substantially. The global average attribution vector lies outside the high-density regions of the explanation manifold (Fig. 6), indicating that averaging explanations across stochastic runs again produces a representation that does not correspond to a typical deployed model instance. Together, these results show that explanation multiplicity persists in Random Forests despite ensemble averaging and high predictive accuracy.
Figure 6: Breast Cancer - Random Forest: universal explanation manifold. PCA embedding of explanation (SHAP) vectors obtained by stacking Random Forest models across all train–test splits. Each point corresponds to a trained ensemble instance and is coloured by predictive accuracy. Multiple dense and separated regions are visible in explanation space, indicating persistent explanation multiplicity across resampled datasets. The black cross denotes the global average attribution vector, computed as the mean of all SHAP vectors across runs, which lies outside the high-density regions of the manifold.
Table 5: Cluster-level summary for Random Forest on the Breast Cancer dataset under stochastic training with varied hyperparameters. Each row corresponds to one explanation cluster within a train-test split. We report cluster size and predictive accuracy (mean $\pm$ std across runs assigned to the cluster). Across splits, two explanation clusters are consistently observed, with comparable predictive performance and stable relative sizes.
Split Cl. Runs Acc. mean Acc. std
100 0 64168 0.967 0.005
100 1 127832 0.974 0.005
101 0 64104 0.956 0.008
101 1 127896 0.959 0.006
102 0 64031 0.953 0.011
102 1 127969 0.956 0.005
103 0 64120 0.946 0.012
103 1 127880 0.949 0.007
104 0 64145 0.941 0.007
104 1 127855 0.945 0.007
Figure 7: Random Forest explanation landscape for Breast Cancer (single split). PCA projection of explanation vectors for split 103 shows two distinct, high-density regions within a single train–test split. The first principal component explains 75.8% of the variance, while PC2 captures residual spread. Colour indicates test accuracy, which varies smoothly across both regions. The persistence of multiple explanation clusters within a single split, and at comparable accuracy, indicates that mechanistic multiplicity in Random Forests is intrinsic and does not rely on aggregation across splits or hyperparameter overlay.
Preliminary experiments on additional datasets and model classes (including Adult Income and Elastic Nets) suggest similar patterns of explanation multiplicity. These analyses are not reported here and are provided in the accompanying code repository as exploratory results.
8 Discussion
EvoXplain shows that even models with high prediction accuracy do not guarantee explanation stability, even when datasets are thought to be ”simple”. Across both datasets, multiple model classes resulted in multiple explanatory modes with considerable support. This suggests that interpretability cannot be inferred from model form or accuracy alone. For example, linear models commonly assumed to be transparent do not consistently converge to a single explanatory mechanism under standard retraining procedures. In our experiments, Logistic Regression exhibits explanation multiplicity across datasets when regularisation is allowed to vary, despite high and stable predictive accuracy (lipton2018myth; rudin2019stop). This challenges the common practice of treating transparency and accuracy as sufficient conditions for interpretability. The goal is not to eliminate variability, but to characterise the range of mechanisms that arise under standard retraining procedures commonly used in practice.
A dominant operational assumption in applied machine learning is that inspecting a single trained model instance is sufficient to characterise a system’s reasoning. EvoXplain clarifies the limits of that assumption, as it shows that models with the same accuracy can have different internal reasoning structures. This implies that one attribution profile is likely just one out of many possible ones. Accordingly, interpretability should be framed as a population-level property of trained models rather than as an attribute of a single fitted instance. While predictive performance remains stable, explanation multiplicity matters because explanations are often used to justify, audit, or regulate model-driven decisions at the individual level. This suggests that, rather than just focusing on a single model instance, we need to develop stability-aware interpretability of models that provides explanations based on the behaviour of models over multiple instances or runs (arrieta2020explainable; selbst2018intuitive).
EvoXplain does not claim to identify where multiplicity originates—whether from data structure, model flexibility, or optimisation randomness. Instead, it makes this uncertainty visible. The framework is deliberately agnostic to causal origin and instead focuses on empirical detectability.
The multiplicity observed cannot be attributed to instability or noise in SHAP itself. SHAP values are deterministic for any fixed model. EvoXplain examines attribution profiles across independent model trainings, rather than variability within a single explanation. This distinction is critical: EvoXplain characterises variability across trained models, not variability within an attribution method. Because SHAP is deterministic for a fixed model, any observed clustering reflects differences between trained models rather than stochasticity in the explainer itself. As such, the resulting clusters reflect real differences in learned mechanisms. For the Breast Cancer benchmark, the centroids of the discovered basins reinforce established diagnostic attributes such as perimeter, radius, area, concavity, and associated morphological descriptors used in clinical tumour assessment. The emergence of distinct, clinically meaningful feature profiles across basins would not be expected if clustering were driven by random noise.
Therefore, inconsistency between explanations is not something to be averaged out as noise. Instead, we reframe it as an empirical observation. In this sense, explanation multiplicity is a property to be measured, reported, and reasoned about, rather than suppressed. In the presence of multiple modes of explanation, the decision to select one becomes a question of science or governance, rather than something the training pipeline can resolve automatically (ecegai2023aiact; rudin2019stop).
This is particularly important because explanations have immediate consequences in certain domains. In criminal justice contexts, multiple basins of attraction counteract the assumption that a model provides a singular, stable justification umbach2023predictive. In biomedicine, instability observed even on standard datasets warns that attribution-based interpretations should be treated with caution when used to support clinical claims agrawal2023breast. Thus, EvoXplain helps to identify cases where a single explanation is insufficient to account for the behaviour of a model under retraining. In both settings, explanation stability becomes relevant precisely because explanations are used as part of decision-making processes rather than as post-hoc descriptions.
In our experiments, we focus on tabular learning, where repeated training is manageable. However, retraining large language models (LLMs) presents a more difficult challenge. We do not claim that the specific geometric structures observed here generalise directly to large-scale models; rather, we argue that the underlying issue of mechanistic non-identifiability is likely to be amplified in such settings. LLMs train in very high-dimensional, non-convex spaces, are exposed to stochastic data, and follow fine-tuning paths that may diverge substantially. Whether models that exhibit similar external behaviour in such settings rely on the same internal mechanisms remains an open empirical question. Measuring this at scale is challenging, but not impossible: models and their fine-tuned variants can be compared, checkpoints sampled, and attribution patterns analysed across independent runs (olah2020circuits). EvoXplain provides a conceptual starting point. The instability observed even in small models suggests that non-identifiability in foundation models may be important and deserves further study (lakshminarayanan2017simple).
Beyond the empirical evidence, mechanistic multiplicity underscores the need to evaluate reliability and interpretability as properties that cannot be assumed. Instead, they must be empirically established under realistic training variability.
Why Mechanistic Multiplicity Matters
Mechanistic multiplicity primarily raises questions of reliability, rather than philosophy. It concerns whether a modelling pipeline produces consistent reasoning under retraining, not whether a single explanation appears plausible. When predictive accuracy decays, retraining is standard practice, and the modelling process may generate different reasoning outputs, with no guarantee of stable reasoning over time. Reliability must therefore concern the process rather than a particular model instance. In high-stakes contexts, explanations justify actions. Even at equal predictive accuracy, models can diverge in the actions, risk factors, and scientific hypotheses they imply, settling into different basins of explanation. This divergence is not resolved by appeals to sparsity, simplicity, or accuracy alone. When explanations inform decisions rather than merely describe them, instability in the underlying mechanism becomes a substantive safety and interpretability issue. Appeals to simplicity do not resolve multiplicity. In the presence of competing explanation basins with equivalent sparsity, complexity, and predictive accuracy, Occam’s razor cannot favour one over another. What remains is a form of empirical underdetermination. In many cases, competing explanation basins are comparable not only in accuracy, but also in sparsity and apparent simplicity, leaving no principled criterion for selection based solely on model-internal properties. In such cases, the presence of multiple stable explanatory modes evidences a profound underdetermination, which EvoXplain is designed to identify and document.
There are boundaries to this work. We analyse explanation stability using attribution methods; other interpretability approaches, such as causal abstraction or circuit tracing, may offer different perspectives (olah2020circuits). EvoXplain does not attempt to correct bias, determine which basin is “correct”, or quantify downstream risks such as fairness or failure modes selbst2018intuitive. We consider only a limited number of datasets and model classes; this study represents an initial step toward broader analysis across architectures, attribution schemes, and domains such as reinforcement learning or language modelling. These limitations define the scope of the present study rather than its validity. The contribution of EvoXplain is not to propose a new explainer, but to provide a diagnostic framework that makes explanation instability measurable, comparable, and reportable across model populations.
Despite these limitations, EvoXplain opens new research questions. Are explanation basins correlated with latent subgroups, causal mechanisms, or inductive bias pathways (marx2020predictive)? Is it possible to prefer certain basins for safety, fairness, or generalisability (rudin2019stop)? More generally, EvoXplain shifts interpretability from analysing individual models to analysing populations of models trained under the same procedure. This reframing allows explanation stability to be measured, compared, and monitored at the level of model populations.
To summarise, EvoXplain treats interpretability as a property of model behaviour under retraining. By focusing on mechanistic multiplicity rather than assuming a single explanatory mode, EvoXplain provides tools to uncover and characterise multiplicity when it arises. This, in turn, raises methodological, epistemic, and governance challenges. If models admit multiple internally consistent mechanistic explanations, then explanation selection, validation, and governance require explicit criteria. If models embody a multiplicity of mechanistic stories, how should they be assessed, described, or governed? While EvoXplain does not resolve these questions, it makes them explicit and positions explanation stability as a concrete target for future research. In practice, EvoXplain enables practitioners and auditors to detect when single-model explanations are insufficient, without prescribing which explanation should be preferred.
9 Conclusion
Explanations are often treated as fixed model properties—examined once and assumed to reflect one underlying mechanism. EvoXplain argues the opposite, and instead quantifies this assumed stability. Across two benchmarks and two model families, models achieved high predictive performance. However, their explanations frequently diverged into multiple distinct modes, even within the same model class and data split. The extent and structure of this multiplicity varied by dataset—for example, explanation basins were more evenly balanced on the Breast Cancer dataset and more asymmetric on COMPAS—but in all cases, high predictive accuracy did not imply a single stable mechanism. This demonstrates that explanation stability cannot be inferred from predictive performance, model class, or perceived simplicity alone.
EvoXplain offers a diagnostic view rather than a decision rule. It shows that multiplicity exists, shows its form, and specifies the frequency at which different explanatory mechanisms can emerge. It does not dictate which local minima should be prioritized; that is a decision to be made from the domain knowledge, validation or regulatory governance. It is a valuable tool to make multiplicity non-abstract and measurable, and to encourage interpretability practices that account for stability. In this sense, EvoXplain complements existing interpretability methods by exposing when single-model explanations are insufficient.
Future research will focus on exploring if mechanistic multiplicity is present within other paradigms and methods of training and attribution. The phenomenon is likely to be much broader than what is presented here, especially in areas where the decisions of the models have real consequences, like genomics and biomedical risk modelling. The question is further amplified within foundation models. Do models that function similarly, internally follow the same mechanism, or do they split into distinct explanatory modes at scale? Answering this question will require population-level analysis of trained models rather than inspection of isolated instances.
Interpretability can no longer centre on a single trained model. The assumed single explanation is viewed as a distribution of explanations, and their stability, diversity, and governance relevance are not assumptions, but empirical matters that require measurement. EvoXplain pushes the field in this direction by treating explanation stability as a primary diagnostic of model behaviour. By making explanation multiplicity observable and quantifiable, EvoXplain provides a foundation for future interpretability, auditing, and governance frameworks that operate at the level of model populations rather than individual fits.
Acknowledgements
The EvoXplain framework is the subject of a provisional patent filing. The experiments reported here were run on the University of Hertfordshire High Performance Computing Cluster, whose support is gratefully acknowledged.
Competing Interests
EvoXplain is the subject of a filed provisional patent application. The author declares no other competing interests.